# Reward Model Ensembles Mitigate but do not Eliminate Reward Hacking

### v2

@@ 3 Underspecification in Reward Models
We now analyze alignment strategies that use a single reward model, and demonstrate that reward models are underspecified. First, Table 2 shows the average in-distribution accuracy across the 25 different reward models, together with the standard deviation (which is low in-distribution).
@@
Figure 2: Average reward of the best-of-$n$ output, as judged by: the same reward model used for ranking (self); reward models fine-tuned from the same pretrain as the ranker (same pretrain); reward models fine-tuned from different pretrains from the ranker (diff pretrain). The reward models that do not share a pretrain with the ranker regard the ranker’s preferred outputs as significantly worse.
@@
The story changes, however, when we move to out-of-distribution data. Figure 2 shows the expected reward achieved by BoN as a function of the number of sampled candidates, $n$, for three reward model scales (KL is approximately $\log n-\frac{n-1}{n}$). The dotted green line shows the expected reward of the top-ranked output according to the reranker itself, while the dashed orange line shows the expected reward of the same output according to reward models that share a pretrain seed. The solid blue line shows the expected reward according to reward models that do not share a pretrain seed. Unsurprisingly, the reranker scores its own top outputs more favorably than the other reward models do. However, the reranker’s outputs are scored significantly less favorably by reward models which do not share a pretrain with the ranker. Reward models that share a pretrain seed with the ranker model overestimate the true reward of the top-ranked output—suggesting that finetune ensembles are not sufficiently diverse because of the shared pretraining state of each of the ensemble’s members. Notably, this gap does not disappear with scale, and is present for base, large, and XL models.
@@
Moving to alignment, differences in estimated rewards induce different policies from the BoN strategy: Figure 3 shows the effects on agreement of the top-ranked summary when reward models do (crosses) or do not (circles) share pretraining seeds. Different reward models tend to produce different 1-best outputs. Again these differences are strongly associated with the pretraining seed: for example, two reward models from different pretrains will choose a different best-of-16 output more than half the time for both tl;dr and helpfulness and in all scales.
Figure 4: Rank correlation of reward scores for tl;dr reward models that share a pretraining seed and models that do not. RLHF alignment increases disagreements between reward models (lower correlation), particularly at low values of $\lambda$ and for reward models that do not share a pretrain.
Last, Figure 4 analyzes the evolution of agreement of the estimated reward scores when performing RLHF on tl;dr for reward models of various scales. Specifically, we align a policy using a single reward model, and then measure how well pairs of reward models agree on the ranking of samples from that policy using Spearman rank correlation. To compute Spearman, we sample 5 completions for each prompt in the validation set from a policy model, at 2K step intervals during RLHF. We compare the agreement between a set of 5 reward models that share the same pre-training seed and a set of 5 that do not (both sets include the reward model used to drive RLHF). For each prompt, we compute Spearman correlation across all ten pairs in each set and report the mean correlation over the pairs. The correlation of models that do not share a pretrain is lower compared to models that share a pretrain seed. Moreover, correlation goes down during RLHF, indicating that the uncertainty about the true reward increases as a result of alignment.
Overall, our analysis demonstrates that (1) different reward models tend to disagree on out-of-distribution data, particularly when the reward models have different pretraining seeds; (2) this propagates to the trained policy model, in the sense that the resulting policy is highly tuned to the preferences of the specific reward model used to drive it; and (3) as a result, the disagreement between reward models tends to increase during alignment. These findings suggest that reward model ensembles might mitigate reward hacking, which we turn to next.
4 Reward Model Ensembles
We describe how to construct reward model ensembles (§4.1), and evaluate their performance (§4.2).
4.1 Pretrain and Finetune Reward Ensembles
We showed that reward models are underspecified—as they are used more in alignment, they induce a stronger distribution shift in the outputs of the policy, which in turns leads to higher disagreement across reward models. Thus, a natural mitigation strategy is to ensemble multiple reward models, under the assumption that different models will have different errors. Aggregating over the scores of the ensemble members will help when some of the ensemble members erroneously assign high reward to a bad output.
Given a set of reward models $\mathcal{M}$, we define the reward of the ensemble to be ${\overline{r}(x,y)=\textrm{agg}(\{r_{m}(x,y)\}_{m\in\mathcal{M}}})$, with agg indicating an aggregation function (Dietterich, 2000; Lakshminarayanan et al., 2017; Raffel et al., 2020; Zaidi et al., 2021). Intuitively, the aggregation function should be conservative, and return a lower score when there is disagreement between the ensemble members. We consider the following simple aggregation function: mean, median, and mean_minus_std, which subtracts the standard deviation of the reward from the mean to penalize high variance. We also experiment with min, but overall find it to be inferior to the alternatives.
We evaluate two types of reward ensembles: pretrain ensembles, where each member was pretrained using a different random seed, [^1] Pretraining does not complete a single epoch over the pretraining data, and thus the data observed by each member of a pretrain ensemble is different (but sampled from the same distribution). and finetune ensembles, where all members share the same pretraining seed, but use a different seed when finetuned on the reward data (which typically includes preference pairs, where one output is preferred over another). In all cases the ensemble contains exactly 5 individual reward models. Pretrain ensembles are significantly more expensive to train, but are more diverse and hence likely to lead to a more robust reward estimate. In fact, Gleave & Irving (2022) reported negative results when using reward ensembles and hypothesized this is due to ensemble members sharing the same underlying pretrained model.
4.2 Experiments
(a) tl;dr
(b) helpfulness
@@ 4.2 Experiments
Figure 5: In best-of-$n$ reranking, pretrain ensemble reward models significantly improve the quality of outputs in the tl;dr summarization task (top) and the helpfulness task, as measured by a T5-XXL model. Full numerical results are in Appendix A.
@@
We now evaluate reward model ensembles across all tasks. Figure 5 shows the results of ensembling in best-of-$n$ reranking, as measured by an XXL-scale fine-tuned reward model. Pretrain ensembles consistently improve performance over individual reward models, especially for higher values of $n$ for both tl;dr and helpfulness. Finetune ensembles, conversely, improve performance in some cases and are comparable in others. For example, on tl;dr a pretrain ensemble with the mean aggregator achieves a win rate of 90% over the SFT outputs at the XL scale, while the win rate of a finetune ensemble with the same mean aggregator is 87.3%. The win rate of the average individual XL-scale reward model is 85.3% (see Table 7). For visual clarity, in Figure 5 we show only two aggregators: mean and mean_minus_std; see Appendix A for results with other aggregators. In general, the differences between aggregators are small, with mean usually performing at, or near, the top. More conservative aggregators (min and mean_minus_std) come out slightly ahead of mean at the smaller scales on tl;dr, suggesting that high variance may be a bigger issue in this setting.
@@
Figure 6 shows the KL-reward trade-off of ensemble reward models in RLHF for tl;dr and helpfulness (evaluated with the finetuned T5-XXL model). In such plots, a better model is one that improves reward and/or reduces the value of KL from the original SFT policy (Gao et al., 2023; Coste et al., 2023). Indeed, similar to BoN, pretrain ensembles consistently outperform both finetune ensembles as well as the average individual model. We present results for the median and mean aggregators for visual clarity, and report full numerical results in Appendix B. In RLHF, KL values are much higher than BoN (which is bounded by $\approx 3.17$ for $n=64$). Consequently, in this setting we witness explicit reward hacking, in which the T5-XXL rewards decrease even as the RLHF objective improves. This happens most prominently for individual models, in many cases for finetune ensembles, and most rarely for pretrain ensembles—where T5-XXL reward scores decrease only when RLHF uses a T5-Base reward model. Thus, our experiments on real data yield more negative conclusions than Coste et al. (2023) about the potential of ensembles to eliminate reward overoptimization.
(a) tl;dr
(b) helpfulness
@@
Figure 6: In RLHF, pretrain ensemble reward models lead to significantly more favorable reward-KL tradeoffs, as judged by a T5-XXL autoeval model. Each point corresponds to training of models to convergence with a particular value of $\lambda$. We show the median aggregator here, full numerical results are in Appendix B.
@@
(a) BoN + tl;dr
(b) BoN + helpfulness
(c) RLHF + tl;dr
(d) RLHF + helpfulness
Figure 7: Using a prompted autoevaluator (PaLM-2-FLAN), ensemble reward models offer significantly better win rates on both tl;dr and helpfulness. Here all reward models are XL-scale.
Figure 8 shows RLHF results for xsum/nli. Here we see a relatively small improvement for ensembles compared to individual models, and a very small difference between pretrain and finetune ensembles. We conjecture this is because xsum/nli optimizes for a particular aspect of the response, namely its factuality. This allows all models to find simple and similar strategies that lead to high reward (for example, emitting short responses with limited content), and thus ensembling does not lead to large gains in performance. We further elaborate on this when discussing limitations of ensembles in §5.
@@ 5 When do Reward Model Ensembles Fail?
We saw that ensembles improve performance according to automatic evaluation metrics. We now conduct a complementary analysis that illustrates that, for some types of errors, ensembling is ineffective. When all reward models share a similar error pattern, this error propagates to the ensemble. Systematic errors across ensemble members can arise due to biases in the finite reward model training data.
@@
To demonstrate this, we manually analyze ensemble outputs to detect frequent errors, and then perform a qualitative analysis. Figure 9 shows the results of this analysis on all three benchmarks. The x-axis corresponds to outputs of the model after training for a certain number of steps, and the y-axis is a statistic of interest (e.g., average output length). We plot the statistic value for the pretrained ensemble (using mean as a representative aggregation function) and for its members. In addition, for tl;dr and helpfulness, where the reward model is trained on the preference data, we show the statistic value on the preference data validation set, conditioned on the label ‘Preferred’ or ‘Rejected’.
@@
  • For helpfulness (Figure 8(a)), outputs tend to be in a format of a list, and thus we write a regular expression that captures this format. The fraction of outputs that have this pattern increases to roughly 50% for 3 members of the ensemble and to the ensemble itself. Looking at the preference data, we do not detect a tendency to produce list outputs in the preferred responses, as the fraction of outputs that matches this format is roughly 8% for both the preferred and rejected responses.
@@
  • For tl;dr (Figure 8(b)), RLHF alignment leads to longer summaries (Singhal et al., 2023) and also outputs that are more extractive, i.e., copy more from the input. Summary length in characters grows substantially for the ensemble and all its members, where for the ensemble, length increases by a factor of two. On the preference data, indeed preferred responses are slightly longer than rejected responses, but much shorter than outputs post-RLHF. We also compute the longest common subsequence (in characters) between the document and the summary and find that it increases for the ensemble from 28.2 to 49.1. Again, the tendency for copying from the document already occurs in the preference data to a small degree, but is amplified by RLHF. [^2] The distribution of outputs in the preference data is not identical to the distribution of outputs before RLHF, and therefore the statistics after zero training steps do not necessarily match those of the preference data.
  • For xsum/nli (Figure 8(c)), training for factuality tends to make summaries shorter. Additionally, precise numbers are typically omitted from the summaries. Figure 9 shows how all members of the ensemble and the ensemble itself exhibit this phenomenon, with length in characters decreasing rapidly, as well as the fraction of examples that contain any numeric value whatsoever.
@@
Overall, these qualitative findings are symptoms of the tendency for different pretrain reward models to learn to associate certain features with high reward. Policy models can then exploit this association, and use these features to produce outputs that are dramatically different from the reward training data, and that achieve (spuriously) high reward for both single reward models and the ensemble.
@@
Why does this happen for both single reward models and reward model ensembles? As one indication, Lakshminarayanan et al. (2017) have proposed distance-awareness, i.e., the ability to quantify the distance of an example from the training set, as a necessary condition for achieving good uncertainty estimates. They showed in a synthetic binary classfication setup that deep ensembles provide good estimates when examples are on the decision boundary, but underestimate uncertainty in areas that are far from the training distribution. In LM alignment, the policy can shift the output distribution away from the decision boundary to areas where all reward models erroneously extrapolate in the same manner. While we focus on ensembles in this work, we hypothesize that the same phenomenon will occur in other approaches for uncertainty estimation that are not distance-aware, such as Monte-Carlo Dropout (Gal & Ghahramani, 2016) and Epistemic Neural Networks (Osband et al., 2021).
@@ 6 Conclusion
In this work, we investigate reward model ensembles as a method for mitigating reward hacking. We find that diversity of the reward ensemble is crucial, and that a pretrain ensemble that contains members that do not share a pretrain seed leads to stronger generalization during alignment when compared to an ensemble whose members share a pretrain seed. However, reward ensembles are not always effective—for example, we find that they can still assign reward based on spurious correlations between the input and the label. If all members of the ensemble capture the same correlations, the ensemble will inherit the same undesirable behaviour. In such cases, the policy can exploit this vulnerability and shift the distribution towards outputs that overuse this correlation, which results in reward hacking. Consequently, reward model ensembles mitigate, but do not fully eliminate, reward hacking. Future work should examine methods for uncertainty estimation that are more robust to the type of distribution shift that occurs during alignment, particularly those that are aware of how different model policy outputs are from the preference data—such as Gaussian processes (Kuss & Rasmussen, 2003; Chu & Ghahramani, 2005; Liu et al., 2020) and conformal prediction under covariate shift (Tibshirani et al., 2019).

### v3

@@ 3 Underspecification in Reward Models
Figure 2: Average reward of the best-of-$n$ output, as judged by: the same RM used for ranking (self); RMs fine-tuned from the same pretrain as the ranker (same pretrain); RMs fine-tuned from different pretrains from the ranker (diff pretrain). The RMs that do not share a pretrain with the ranker regard the ranker’s preferred outputs as significantly worse.
@@
We begin by demonstrating that the out-of-distribution performance of individual RMs is underspecified. Table 1 shows the average in-distribution accuracy across the 25 different RMs, together with the standard deviation. The low standard deviation highlights that all 25 reward models achieve similar performance in-distribution. But when we move to out-of-distribution data, the models diverge. Figure 2 shows the expected reward achieved by BoN as a function of the number of sampled candidates, $n$. The dotted green line shows the expected reward of the top-ranked output according to the reranker itself, while the dashed blue line shows the expected reward of the same output according to RMs that share a pretrain seed. The solid orange line shows the expected reward according to RMs that do not share a pretrain seed. Unsurprisingly, the reranker scores its own top outputs more favorably than the other RMs do. However, the reranker’s outputs are scored significantly less favorably by RMs which do not share a pretrain with the ranker. RMs that share a pretrain seed with the ranker model overestimate the true reward of the top-ranked output—suggesting that finetune ensembles are not sufficiently diverse because of the shared pretraining state of each of the ensemble’s members. Notably, this gap does not disappear with scale, as shown in Figure 2.
Moving to alignment, differences in estimated rewards induce different policies from the BoN strategy. Figure 8 (in the appendix) shows the effects on agreement of the top-ranked summary when RMs do (crosses) or do not (circles) share pretraining seeds. Different RMs tend to produce different 1-best outputs, and these differences are linked to the pretraining seed: for example, two RMs from different pretrains will choose a different best-of-16 output more than half the time for both tl;dr and helpfulness and in all scales.
Figure 3: Rank correlation of reward scores for tl;dr RMs that share a pretraining seed and models that do not. RLHF alignment increases disagreements between RMs (lower correlation), particularly at low values of $\lambda$ and for RMs that do not share a pretrain.
Figure 3 analyzes the evolution of agreement of the estimated reward scores when performing RLHF on tl;dr. For each policy model, we sample five completions for each prompt in the validation set at 2000 step intervals during RLHF. We then measure how well pairs of RMs agree on the ranking of these completions, using Spearman rank correlation. The correlation is averaged across all pairs of reward models that do and do not share the same pre-training seed; both sets of RMs include the one used to drive RLHF. As shown in the figure, RLHF decreases the average correlation, particularly at low levels of regularization. Furthermore, agreement is substantially lower for pairs of reward models that do not share the same pretraining seed. This supports our conclusion that reward modeling under alignment-induced distribution shift is underspecified by the preference annotations.
Overall, our analysis demonstrates that (1) different RMs tend to disagree on out-of-distribution data, particularly when the RMs have different pretraining seeds; (2) this propagates to the trained policy model, in the sense that the resulting policy is highly tuned to the preferences of the specific RM used to drive it; and (3) as a result, the disagreement between RMs tends to increase during alignment. These findings suggest that reward model ensembles might mitigate reward hacking, which we turn to next.
4 Reward Model Ensembles
A natural mitigation to reward model underspecification is to ensemble multiple RMs, under the assumption that different models will have different errors. Aggregating over the scores of the ensemble members will help when some of the ensemble members erroneously assign high reward to a bad output.
Building reward model ensembles
Given a set of RMs $\mathcal{M}$, we define the reward of the ensemble to be ${\overline{r}(x,y)=\textrm{agg}(\{r_{m}(x,y)\}_{m\in\mathcal{M}}})$, with agg indicating an aggregation function (Dietterich, 2000; Lakshminarayanan et al., 2017; Raffel et al., 2020; Zaidi et al., 2021). Intuitively, the aggregation function should be conservative, and return a lower score when there is disagreement between the ensemble members. We consider the following simple aggregation functions: mean, median, and mean_minus_std, which subtracts the standard deviation of the reward from the mean to penalize high variance. We also experiment with min, but overall find it to be inferior to the alternatives.
We evaluate two types of reward ensembles: pretrain ensembles, where each member is pretrained using a different random seed, [^3] Pretraining does not complete a single epoch over the pretraining data, and thus the data observed by each member of a pretrain ensemble is different (but sampled from the same distribution). and finetune ensembles, where members share the same pretraining seed, but use a different seed when finetuned on the reward data. In all cases the ensemble contains five individual RMs. Pretrain ensembles are expensive to train, but are more diverse and hence likely to lead to a more robust reward estimate. (Gleave & Irving (2022) reported negative results when using reward ensembles and hypothesized this is due to members sharing the same underlying pretrained model.)
@@ 4 Reward Model Ensembles
Figure 4: In BoN reranking, pretrain ensemble RMs significantly improve output quality, as measured by a T5-XXL autoeval model. Full numerical results are in Appendix A.
Evaluation
@@
We now evaluate RM ensembles across tasks. Figure 4 shows the results in best-of-$n$ reranking, as measured by an XXL-scale fine-tuned RM. Pretrain ensembles consistently improve performance over individual RMs, especially for higher values of $n$ for both tl;dr and helpfulness. Finetune ensembles, conversely, improve performance in some cases and are comparable in others. For example, on tl;dr a pretrain ensemble with the mean aggregator achieves a win rate of 90% over the SFT outputs at the XL scale, while the win rate of a finetune ensemble with the same aggregator is 87.3%. The win rate of the average individual XL-scale RM is 85.3% (see Table 6). For visual clarity, in Figure 4 we show only two aggregators: mean and mean_minus_std; see Appendix A for the other aggregators. In general, differences between aggregators are small, with mean usually performing at, or near, the top. More conservative aggregators (min and mean_minus_std) come out slightly ahead of mean at the smaller scales on tl;dr, suggesting that high variance may be a bigger issue in this setting.
@@
Figure 5 shows the KL-reward trade-off of ensemble RMs in RLHF for tl;dr and helpfulness (evaluated with the finetuned T5-XXL model). In such plots, a better model improves reward and/or reduces KL from the SFT policy (Gao et al., 2023; Coste et al., 2023). As with BoN alignment, pretrain ensembles consistently outperform finetune ensembles as well as the average individual RM. For clarity, the figure presents results only for the median aggregator for clarity, but these conclusions are supported by full numerical results across aggregators, shown in Appendix B. RLHF leads to much higher KL values than BoN. Consequently, we witness explicit reward hacking, in which the T5-XXL rewards decrease even as the RLHF objective improves. This happens most prominently for individual models, in many cases for finetune ensembles, and most rarely for pretrain ensembles—where T5-XXL reward scores decrease only when RLHF uses a T5-Base RM. Thus, our experiments on real data yield more negative conclusions than Coste et al. (2023) about the potential of ensembles to eliminate reward overoptimization.
@@
Figure 5: In RLHF, pretrain ensemble RMs lead to more favorable reward-KL tradeoffs, as judged by a T5-XXL autoeval model. Each point corresponds to training to convergence at a particular value of $\lambda$. We show median ensembles here; for others see Appendix B.
@@
(a) BoN
(b) RLHF
Figure 6: Using a prompted autoevaluator (PaLM-2-Flan), ensemble RMs get significantly better win rates on both tl;dr and helpfulness. Here all RMs are XL-scale.
Figure 9 (in the appendix) shows RLHF results for xsum/nli. Here ensembles offer relatively small improvements, and there is little difference between pretrain and finetune ensembles. We conjecture this is because xsum/nli optimizes specifically for factuality. This allows all models to find simple and similar strategies that lead to high reward (namely, emitting short responses), and thus ensembling does not lead to large gains in performance.
@@ 5 When do Reward Model Ensembles Fail?
Ensembles improve performance according to automatic evaluation metrics, but does this mean that reward overoptimization is solved? We now investigate five specific “reward hacks”, and show that they are not prevented by ensembling, because all members agree that these hacks improve the quality of the outputs.
@@
To demonstrate this, we manually identify five qualitative distribution shifts in which the post-RLHF policies differ dramatically from both the reference policy and the preference annotations. Figure 7 and Figure 10 (in the appendix) show the results of this analysis on all benchmarks with a T5-large RM. The x-axis corresponds to the number of RLHF training steps, and the y-axis is a statistic of interest (e.g., average output length). We plot the statistic for the pretrained ensemble (using mean as a representative aggregation function) and for its members. For tl;dr and helpfulness, where the reward model is trained on preference data, we show the statistic value on the preference data validation set, conditioned on the label ‘Preferred’ or ‘Rejected’.
@@
  • For helpfulness (Figure 7(a)), outputs tend to be in a list format, which we capture with a regular expression. The fraction of outputs that have this pattern increases to roughly 50% for three members of the ensemble, and for the ensemble itself. We do not detect this tendency in the preference data: the fraction of outputs that matches this format is roughly 8% for preferred and rejected responses.
@@
  • For tl;dr (Figure 7(b), Figure 10(b)), RLHF leads to longer summaries (Singhal et al., 2023) and more extractive outputs, i.e., more copying from the input. Length in characters grows substantially for the ensemble and its members, where for the ensemble, length increases by a factor of two. On the preference data, preferred responses are slightly longer than rejected responses, but much shorter than outputs post-RLHF. We also compute the longest common subsequence (in characters) between the document and the summary and find that it doubles after RLHF with an ensemble RM. Although preference data shows a slight preference for copying, this preference is dramatically amplified by RLHF.
  • For xsum/nli (Figure 7(c), Figure 10(c)), training for factuality makes summaries shorter and less specific, as measured by omission of numerical values. All members of the ensemble exhibit this phenomenon: RLHF leads to rapid decreases in the length of the outputs and the fraction of outputs that contain numerical values.
(a) helpfulness: fraction of answers containing lists
(b) tl;dr: output length
(c) xsum/nli: fraction of numerical tokens in the output.
Figure 7: Limitations of reward model ensembles. The x-axis is RLHF steps, the y-axis plots different statistics of the average validation output at that step, and the curves correspond to the pretrain ensemble (solid blue) and its members (dashed orange). For preference data, we plot the same statistics conditioned on the preference data label (Preferred vs. Rejected). The statistics of the “aligned” outputs are far from their values in the preference data.
@@
These findings illustrate the tendency of different reward models to associate certain features with high reward. Policy models can exploit this association, using these features to produce outputs that are dramatically different from the reward training data, which achieve (spuriously) high reward for both individual reward models and the ensemble.
@@
Why do reward model ensembles fail to capture uncertainty about these reward hacks? Prior work has shown that ensembles can provide good uncertainty estimates around the decision boundary, while underestimating uncertainty for examples far from the training distribution (Lakshminarayanan et al., 2017). In LM alignment, the policy can shift the output distribution away from the decision boundary to areas where all reward models erroneously extrapolate in the same manner. The same phenomenon may occur in other approaches for uncertainty estimation that are not distance-aware, such as Monte Carlo Dropout (Gal & Ghahramani, 2016) and Epistemic Neural Networks (Osband et al., 2021).
@@ 6 Conclusion
While reward models can improve robustness to alignment-induced distribution shift, diversity of the ensemble is crucial. Pretrain ensembles are more diverse than finetune ensembles, and therefore lead to stronger generalization. However, even pretrain ensembles are not diverse enough: for many reward hacks, all members of the ensemble agree, making the ensemble as vulnerable as its constituents. To summarize, reward model ensembles mitigate, but do not eliminate, reward hacking. Future work should examine uncertainty quantification techniques that are more robust to the type of distribution shift that occurs during alignment, particularly techniques that explicitly represent distributional shift from the preference annotations (Chu & Ghahramani, 2005; Tibshirani et al., 2019).
7 Reproducibility
To support reproducibility and enable further work on pretrain ensembles, we have released all $5\times 3=15$ pretraining checkpoints, covering the t5-base, t5-large, and t5-xl scales. The release can be found at https://github.com/google-deepmind/reward-ensembles. To our knowledge, this is the first release of multiple T5 checkpoints, following prior work on BERT (Sellam et al., 2021). All datasets used in this research are public. Appendix C provides all relevant hyperparameters used during RLHF.